OpenStudy (anonymous):

What is the equation of the ellipse with co-vertices (0, 2), (0, -2) and vertices (3, 0), (-3, 0)?

OpenStudy (anonymous):

Do you know equation of an ellipse?

OpenStudy (anonymous):

no

OpenStudy (freckles):

\[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1 \\ \text{ count right and left } a \text{ from center } \\ \text{ count up and down } b \text{ from center }\]